OpenStudy (lizz123):

HELP 4x + 1 = 2x + 7 CHOICES: a)5 b)3 c)4 d)2

OpenStudy (lizz123):

@chosenmatt

OpenStudy (chosenmatt):

x=3

OpenStudy (anonymous):

you need to try to get all the x terms on one side and the other terms on the other side.

OpenStudy (anonymous):

hey lizz its c

OpenStudy (lizz123):

thnx

OpenStudy (anonymous):

can i get medal

OpenStudy (chosenmatt):

no problem

OpenStudy (zale101):

@chosenmatt we are not allowed to give direct answers, but to help them through it instead.

OpenStudy (anonymous):

4x + 1 = 2x + 7 --- subtract 2x from both sides 4x - 2x + 1 = 7 --- subtract 1 from both sides 4x - 2x = 7 - 1 -- combine like terms 2x = 6 -- divide by 2 x = 3

OpenStudy (chosenmatt):

Simplifying 4x + 1 = 2x + 7 Reorder the terms: 1 + 4x = 2x + 7 Reorder the terms: 1 + 4x = 7 + 2x Solving 1 + 4x = 7 + 2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2x' to each side of the equation. 1 + 4x + -2x = 7 + 2x + -2x Combine like terms: 4x + -2x = 2x 1 + 2x = 7 + 2x + -2x Combine like terms: 2x + -2x = 0 1 + 2x = 7 + 0 1 + 2x = 7 Add '-1' to each side of the equation. 1 + -1 + 2x = 7 + -1 Combine like terms: 1 + -1 = 0 0 + 2x = 7 + -1 2x = 7 + -1 Combine like terms: 7 + -1 = 6 2x = 6 Divide each side by '2'. x = 3 Simplifying x = 3